# Semantic Superiority vs. Forensic Efficiency: A Comparative Analysis of Deep Learning and Psycholinguistics for Business Email Compromise Detection

### v1

@@ Abstract
Business Email Compromise (BEC) is a sophisticated social engineering threat that manipulates organizational hierarchies and exploits psychological vulnerabilities like trust and authority, leading to significant financial and reputational damage for enterprises. According to the 2024 FBI Internet Crime Report, BEC accounts for over \$2.9 billion in annual adjusted losses, presenting a significant economic asymmetry: the cost of a False Negative (fraud loss) exceeds the cost of a False Positive (manual review) by orders of magnitude (approximately 1 to 5,480).
This paper examines two distinct detection paradigms for BEC: the Forensic Psycholinguistic Stream, which utilizes CatBoost to analyze psycholinguistic cues with high interpretability and low latency, and the Semantic Stream, which employs DistilBERT for deep learning-based contextual language understanding, offering superior accuracy at the cost of higher computational resources. We evaluated DistilBERT on an adversarially poisoned dataset (N = 7,990) generated via our Black Hole protocol. It was benchmarked on Tesla T4 GPU infrastructure, achieving superior detection (AUC = 1.0000, F1 = 0.9981) with acceptable real-time latency (7.403 milliseconds). CatBoost achieves competitive detection (AUC = 0.9905, F1 = 0.9486) at 8.4x lower latency (0.885 milliseconds), consuming negligible computational resources. For organizations with GPU infrastructure, DistilBERT offers superior accuracy, making it an ideal choice. On the other hand, CatBoost is preferable for edge deployments or cost-sensitive environments due to its comparable security and lower operational costs, providing a viable alternative. Both approaches demonstrate a return on investment exceeding 99.96% when optimized through cost-sensitive learning, as evidenced by their ability to significantly reduce false negatives and associated financial losses. This high return is achieved by minimizing the financial impact of undetected fraudulent emails, which are substantially more costly than false positives.
@@ I-B Problem Statement
    Economic Asymmetry: The cost of a False Positive ($C_{FP}$) is approximately \$25. The cost of a False Negative ($C_{FN}$) is approximately \$137,000 on average [1]. Standard loss functions fail to capture this 1:5,480 risk ratio.
@@ IV-A Psycholinguistics: Grice’s Maxims & The “Smiling Assassin”
Legitimate business communication typically adheres to Grice’s Maxim of Quality (be truthful). BEC attacks violate this while adhering to the Maxim of Manner (be polite) to mask the deception. We quantify this “Polite Coercion” using the Smiling Assassin Score ($\psi$).
@@ VI-B Modern Data Validation (PMCC-2025)
To test temporal generalization, we curated a private corpus of 1,000 legitimate business emails from 2025 (PMCC-2025), containing modern slang and mobile signatures.
@@ X Conclusion
DistilBERT achieves superior detection (AUC 1.0000, F1 0.9981) with acceptable real-time latency (7.403 milliseconds) when GPU resources are available. CatBoost achieves competitive detection (AUC 0.9905, F1 0.9486) at 8.4x lower latency (0.885 milliseconds), consuming negligible GPU or CPU resources. For organizations with GPU infrastructure (cloud providers, large enterprises), DistilBERT provides maximum accuracy with acceptable latency, while for edge deployments, CPU-only gateways, or cost-sensitive environments, CatBoost delivers near-equivalent security with reduced operational overhead. This work demonstrates that the architecture choice depends on infrastructure availability, as semantic models with GPU acceleration provide maximum accuracy while forensic models are suited for resource-constrained scenarios. Both approaches exceed 99.96% accuracy.
@@ Acknowledgment
The author thanks Dr. Frederick Ayivor, Ph.D, for guidance and support throughout this work. His advice on research design, analysis and interpretation helped shape the direction of the study. The author also thanks the open-source Kaggle community for providing datasets and tools that enabled data processing and initial experiments.

### v2

@@ Abstract
Business Email Compromise (BEC) is a sophisticated social engineering threat that manipulates organizational hierarchies, leading to significant financial damage. According to the 2024 FBI Internet Crime Report, BEC accounts for over \$2.9 billion in annual losses, presenting a massive economic asymmetry: the financial cost of a False Negative (fraud loss) exceeds the operational cost of a False Positive (manual review) by a ratio of approximately 5,480:1.
This paper contrasts two detection paradigms: a Forensic Psycholinguistic Stream (CatBoost), which analyzes linguistic cues like urgency and authority with high interpretability, and a Semantic Stream (DistilBERT), which utilizes deep learning for contextual understanding. We evaluated both streams on a hybrid dataset (N=7,990) containing human-legitimate and AI-synthesized adversarial fraud.
Benchmarked on Tesla T4 infrastructure, DistilBERT achieved near-perfect detection on synthetic threats (AUC >> 0.99, F1 = 0.998) with acceptable real-time latency (7.4 ms). CatBoost achieved competitive detection (AUC = 0.991, F1 = 0.949) at 8.4x lower latency (0.8 ms) with negligible resource consumption. We conclude that while DistilBERT offers maximum accuracy for GPU-equipped organizations, CatBoost provides a viable, cost-effective alternative for edge deployments. Both approaches demonstrate a theoretical ROI exceeding 99.9% when optimized via cost-sensitive learning.
@@ I-B Problem Statement
    Economic Asymmetry: The operational reality is defined by cost, not just frequency. The cost of a False Positive ($C_{FP}$) is approximately \$25, while the cost of a False Negative ($C_{FN}$) is approximately \$137,000 on average [1]. Standard loss functions fail to capture this 1:5,480 risk ratio.
@@ III-B Risk Scenarios (MITRE ATT&CK)
III-C Limitations & Threats to Validity
Synthetic fraud constitutes 52.7% of the dataset. While GPT-4 generation enables adversarial diversity, it may not capture operational evasion tactics employed by sophisticated threat actors. The Enron corpus (2001-2002) introduces temporal bias; linguistic conventions have shifted significantly. PMCC-2025 (N=1,000) partially addresses this by testing on modern business communication patterns, but comprehensive evaluation requires longitudinal real-world deployment data unavailable at this stage.
@@ IV-A Psycholinguistics: Grice’s Maxims & The “Smiling Assassin”
Legitimate business communication typically adheres to Grice’s Maxim of Quality (be truthful). BEC attacks violate this while adhering to the Maxim of Manner (be polite) to mask the deception. We operationalize this dynamic as an engineered feature—the Smiling Assassin Score ($\psi$)—capturing politeness-urgency interaction. While inspired by Gricean pragmatics, this metric requires empirical validation against annotated deception corpora to establish psycholinguistic validity.
@@ VI-B Modern Data Validation (PMCC-2025)
To test temporal generalization, we curated a private corpus of 1,000 legitimate business emails from 2025 (PMCC-2025), containing modern slang and mobile signatures. PMCC-2025 contains contemporary linguistic patterns absent from the Enron-based test set: mobile signatures, emoji usage, informal syntax. This validates temporal generalization models trained on 2001-2002 data must handle 2025 communication styles without catastrophic drift.
@@ IX-B Cost Surface & ROI
The optimal threshold ($\tau_{L}\approx 0$) maximizes financial utility under fixed cost assumptions. However, operational deployment introduces analyst workload constraints and alert fatigue risks not captured in this model. Organizations processing 100K+ emails daily may require higher thresholds to maintain sustainable review queues, trading marginal detection gains for operational feasibility.
@@ X Conclusion
DistilBERT achieves superior detection (AUC 1.0000, F1 0.9981) with acceptable real-time latency (7.403 milliseconds) when GPU resources are available. CatBoost achieves competitive detection (AUC 0.9905, F1 0.9486) at 8.4x lower latency (0.885 milliseconds), consuming negligible GPU or CPU resources. For organizations with GPU infrastructure (cloud providers, large enterprises), DistilBERT provides maximum accuracy with acceptable latency, while for edge deployments, CPU-only gateways, or cost-sensitive environments, CatBoost delivers near-equivalent security with reduced operational overhead. This work demonstrates that the architecture choice depends on infrastructure availability, as semantic models with GPU acceleration provide maximum accuracy while forensic models are suited for resource-constrained scenarios. Both approaches exceed 99.96% accuracy. Future work must validate performance on longitudinal operational data to address synthetic bias and temporal drift. Psycholinguistic feature constructs require empirical validation through controlled deception studies.
@@ Acknowledgment
The authors thank the open-source Kaggle community for providing datasets and tools that enabled data processing and initial experiments.
@@ Appendix A Feature Definitions
Table VII lists the psycholinguistic features used in the analysis. It defines each feature and states what linguistic attribute it captures, including politeness–urgency interaction, authority cues, hedge frequency, sentiment shift, and uppercase character ratio.